equation of a line containing the given point and perpendicular to the given line: (4,-3);2x+7y=8
Found 2 solutions by robertb, amoresroy:
Answer by robertb(5830)

The desired equation would have the form 7x+-+2y+=+c, where c is a constant that is to be determined. To find c, plug in the coordinates of the given point into the desired equation: 7*4 - 2*-3 = c, or 28 + 6 = c, or c = 34. Therefore the eqaution is 7x+-+2y+=+34.

Answer by amoresroy(361)


2x+7y=8
7y = -2x + 8
y = -2/7*x + 8/7
The slope of perpendicular line is negative reciprocal of -2/7
m = 7/2
y = mx + b
solve for b, given x =4, y=-3
-3 = 7/2 * 4 + b
-3 = 14 + b
b = -17
So equation of line is
y = 7/2 * x - 17
2y = 7x - 17
17 = 7x - 2y
7x - 2y = 17